# On the cohomological dimension of Siegel modular varieties and the modularity of formal Siegel modular forms

### v1

@@ 1 Introduction
In Kudla’s famous paper [Kud97], he conjectured the modularity of the generating series of special cycles. There are fruitful results in this direction, for example, assuming absolutely convergence of the generating series, Wei Zhang proved the modularity in the Siegel case in [Zha09] and Yifeng Liu in the unitary case in [Liu12]. To remove the convergence condition, Bruinier and Raum proved the modularity of formal Fourier–Jacobi series in the Siegel case in [BR15] and Jiacheng Xia in some unitary case in [Xia22] by using technical analytic method. However, recently Bruinier and Raum presented a new approach in this automatic convergence problem in the Siegel case in [BR24], which is highly related to my results on the coherent cohomological dimension of Shimura varieties. Actually, in an ongoing project joint with Wenxuan Qi, Linli Shi, Peihang Wu, Liang Xiao and Yichao Zhang, we have a new approach of the modularity conjecture, where the cohomological dimension appears naturally.
@@
We fix a field $k$ that is algebraically closed and characteristic 0. When we talk about variety, we mean a reduced separated scheme of finite type over $k$, but it does not need to be irreducible. Let $A_{g,\Gamma}$ be the Siegel modular variety of genus $g\geq 2$ and arithmetic level $\Gamma\subset\operatorname{Sp}_{2g}(\mathbb{Q})$ (i.e. $\Gamma$ is commensurable with $\operatorname{Sp}_{2g}(\mathbb{Z})$ and contains a principal congruence subgroup $\Gamma(N)$ for some $N$). Our main result is the following.
@@
As a key application, we deduce that formal Siegel modular forms are automatically convergent, i.e., they are Siegel modular forms. This generalizes earlier work of Bruinier and Raum [BR15] and [BR24].
@@
To establish Theorem 1.1, we introduce and study weak G2 and G3 properties with respect to a line bundle, which refine the classical notions of G2 and G3 subvarieties in the sense of Hironaka–Matsumura in [HM68].
@@
Let $A^{\min}_{g,\Gamma}$ denote the minimal compactification, and let $A^{\Sigma}_{g,\Gamma}$ denote the toroidal compactification with respect to the cone decomposition $\Sigma$. Let $\omega$ denote the Hodge line bundle on $A_{g,\Gamma}^{\min}$ or its pullback on $A_{g,\Gamma}^{\Sigma}$. We use $D_{\Gamma}^{\min}$ and $D_{\Gamma}^{\Sigma}$ to denote the corresponding boundary of the compactifications. There is a set-theoric description of $A_{g,\Gamma}^{\min}$ that
@@
and $D_{\Gamma,l}^{\Sigma}$ be its preimage in $A_{g,\Gamma}^{\Sigma}$.
@@
Acknowledgement. I would like to thank Liang Xiao for asking me to consider this problem. And I want to thank Jan Hendrik Bruinier, Wenxuan Qi, Martin Raum, Yichao Tian, Zhiyu Tian, Peihang Wu, Deding Yang, Zhiwei Yun, Wei Zhang, and Yichao Zhang for helpful discussions.
@@ 2 Properties G2 and G3
By [Sta25, Tag 03GE], taking integral closure commutes with étale base change, so the considered diagram is Cartesian. ∎
@@ 3 Weak G2 and G3 properies with respect to a line bundle
It is obvious that $K(X,\mathcal{L})$ is independent of the choice of the embedding.
@@
The part $(1)$ is trivial. For $(2)$, it is trivial that $K(X,\mathcal{L}^{n})\subset K(X,\mathcal{L})$. Conversely, if $s_{0}/s_{1}\in K(\mathcal{L})$ such that $s_{0},s_{1}\in H^{0}(X,\mathcal{L}^{m})$, then $s_{0}s_{1}^{n-1},s_{1}^{n}\in H^{0}(X,\mathcal{L}^{nm})$ hence $s_{0}/s_{1}=s_{0}s_{1}^{n-1}/s_{1}^{n}$ lies in $K(X,\mathcal{L}^{n})$.
For $(3)$, using $(2)$, we may assume that $\mathcal{L}$ is very ample. This gives a closed immersion $j:X\hookrightarrow\mathbb{P}^{N}$ over $k$ such that $j^{*}\mathcal{O}_{\mathbb{P^{N}}}(1)=\mathcal{L}$, and the coordinates are given by $s_{0},\dots,s_{N}\in H^{0}(X,\mathcal{L})$. Then $K(X)$ can be generated by ratios of $s_{0},\dots,s_{N}$ hence is equal to $K(X,\mathcal{L})$. ∎
@@
be the natural quotient morphism. Let $Y$ be a closed subvariety of $X$ such that both $K(X_{/Y})$ and $K(X^{\prime}_{/f^{-1}(Y)})$ are fields. Let $\mathcal{L}$ be a line bundle on $X$ and $\mathcal{L^{\prime}}=f^{*}\mathcal{L}$, such that $K(X,\mathcal{L})=K(X)$ and $K(X^{\prime},\mathcal{L^{\prime}})=K(X^{\prime})$. Then $Y$ is weakly G2 (resp. G3) in $X$ with repect to $\mathcal{L}$ if and only if $f^{-1}(Y)$ is G2 (resp. G3) in $X^{\prime}$ with repect to $\mathcal{L}^{\prime}$.
@@
It’s obvious that $G$ is the Galois group of $K(X^{\prime})$ over $K(X)$. The action of $G$ on $X^{\prime}$ induces a natural group action on $H^{0}(X^{\prime}_{/f^{-1}(Y)},\hat{\mathcal{L}^{\prime}})$ and we have
@@
Let $f:X^{\prime}\to X$ be a proper birational morphism of projective varieties, and $Y$ be a closed subvariety of $X$ such that both $K(X_{/Y})$ and $K(X^{\prime}_{/f^{-1}(Y)})$ are fields. Assume further that $X$ is normal. Let $\mathcal{L}$ be a line bundle on $X$ and $\mathcal{L}^{\prime}=f^{*}\mathcal{L}$. Then $Y$ is weakly G2 (resp. G3) in $X$ with repect to $\mathcal{L}$ if and only if $f^{-1}(Y)$ is G2 (resp. G3) in $X^{\prime}$ with repect to $\mathcal{L}^{\prime}$.
@@
It’s sufficient to show that $H^{0}(X^{\prime}_{/f^{-1}(Y)},\hat{\mathcal{L}}^{\prime\otimes n})=H^{0}(X_{/Y},\hat{\mathcal{L}}^{\otimes n})$. We show that $f_{*}\mathcal{O}_{X_{/f^{-1}(Y)}^{\prime}}=\mathcal{O}_{X_{/Y}}$ then above equality holds by projection formula. By Zariski’s main theorem we have $f_{*}\mathcal{O}_{X^{\prime}}=\mathcal{O}_{X}$, thus we have $f_{*}\mathcal{O}_{X_{/f^{-1}(Y)}^{\prime}}=\mathcal{O}_{X_{/Y}}$ by flatness of $X_{/Y}\to X$. ∎
@@
Let $g:X_{2}\to X_{1}$ be a proper morphism of normal projective varieties, and let $\mathcal{L}_{1}$ be a line bundle on $X_{1}$ and $\mathcal{L}_{2}=g^{*}\mathcal{L}_{1}$. Let $Y_{1}$ be a closed subvariety in $X_{1}$ and $Y_{2}=f^{-1}(Y_{1})$. Assume that $Y_{1}$ is weakly G2 with repect to $\mathcal{L}_{1}$ and $Y_{2}$ is weakly G2 with repect to $\mathcal{L}_{2}$ in $X_{2}$.
@@
The proof is same as Proposition 2.12 by using commutative diagram
@@ 4 Cohomological dimension and Grothendieck–Lefschetz condition
In this section, we review the notion of (coherent) cohomological dimension and Grothendieck–Lefschetz condition, and how they related to the weak G2 and G3 properties.
@@
The following lemma relates $\operatorname{ccd}(X)$ to vanishing of cohomology of negative powers of an ample line bundle.
@@
Let $\mathcal{O}_{X}(1)$ be an ample line bundle on $X$. Write $\mathcal{O}_{X}(m)$ for its $m$-th power. Then $\operatorname{ccd}(X)\leq r$ if and only if there exists $M\in\mathbb{Z}_{\geq 0}$ that $H^{i}(X,\mathcal{O}_{X}(-m))=0$ for all $i>r$ and $m\geq M$.
@@
It suffices to prove the if part. We prove $\operatorname{ccd}(X)\leq i$ by reverse induction on $i\geq r$. When $i>\dim X$, this is trivial. Suppose that $\operatorname{ccd}(X)\leq i+1$ and $i\geq r$, we prove that $\operatorname{ccd}(X)\leq i$. For any coherent sheaf $\mathcal{F}$ on $X$, we have $\mathcal{F}(m):=\mathcal{F}\otimes\mathcal{O}_{X}(m)$ is generated by global sections for sufficient large $m$ (and $m\geq M$). Then we have an exact sequence
$0\longrightarrow\mathcal{G}\longrightarrow\mathcal{O}_{X}^{N}\longrightarrow\mathcal{F}(m)\longrightarrow 0$
with coherent sheaf $\mathcal{G}$, and then exact sequence
$0\longrightarrow\mathcal{G}(-m)\longrightarrow\mathcal{O}_{X}^{N}(-m)\longrightarrow\mathcal{F}\longrightarrow 0.$
@@
$0=H^{i}(X,\mathcal{O}_{X}^{N}(-m))\longrightarrow H^{i}(X,\mathcal{F})\longrightarrow H^{i+1}(X,\mathcal{G}(-m))=0$
@@
Now by Serre duality, we have
@@
For local cohomology we have the exact sequence
@@
and by the condition $\operatorname{ccd}(X\backslash Y)\leq d-2$ we have
@@
Comining the duality results above, we have
@@
As $\mathcal{E}_{U}$ is locally free, we have an injective sheaf homomorphism $\mathcal{E}_{U}\hookrightarrow\hat{i}_{*}\hat{i}^{*}\mathcal{E}_{U}$ on $U$, and it implies injections
@@
$(3)\Rightarrow(1)$: Since $\mathcal{O}(1)$ is an ample line bundle on $X$, it is sufficient to prove that $H^{i}(X\backslash Y,\mathcal{O}(-m))=0$ when i>d2i>d-2 and $m$ large enough by Lemma 4.2. Again we use the condition $\mathrm{Lef}(X,Y)$ and the duality theorems above, we have
$H^{d}_{Y}(X,\mathcal{O}(-m))\cong H^{d}(X,\mathcal{O}(-m)).$
@@
$H^{d-1}(X,\mathcal{O}(-m))\cong H^{1}(X,\omega_{X}\otimes\mathcal{O}(m))=0$
for $m$ large enough. Now the expected vanishing follows from the exact sequence
$0\to H^{d-1}(X\backslash Y,\mathcal{O}(-m))\to H^{d}_{Y}(X,\mathcal{O}(-m))\xrightarrow{\sim}H^{d}(X,\mathcal{O}(-m))\to H^{d}(X\backslash Y,\mathcal{O}(-m))\to 0.$
The following theorem of Hartshorne and Speiser states a relation between the G3 property and the Grothendieck–Lefschetz condition. See in [Bad04, Thm. 10.7] or [Har06, Chap.V, Prop. 2.1].
@@
We need a slightly different version which replaces G3 with weakly G3 with repect to an ample line bundle.
@@
Let $Y$ be a closed subvariety of a projective variety $X$ (of dimension $d\geq 2$) that locally satisfies Serre’s condition $S_{2}$ (e.g. when $X$ is normal). Assume that $Y$ is weakly G3 in $X$ with respect to an ample line bundle $\mathcal{L}$, and $Y$ intersects nontrivially with every effective divisor of $X$. Then $(X,Y)$ satisfies the condition $\mathrm{Lef}(X,Y)$.
@@
The proof is almost the same as the proof of the above theorem. For an integer $m\geq 0$, we fix an embedding $\mathcal{L}^{m}\hookrightarrow K_{X}$, and then it induces an embedding $\hat{\mathcal{L}}^{m}:=\hat{i}^{*}\mathcal{L}^{m}\to\mathcal{M}_{X_{/Y}}$ where $\hat{i}:X_{/Y}\to X$ is the natural morphism. We have the commutative diagram
@@
Now we prove that $(X,Y)$ satisfies $\mathrm{Lef}(X,Y)$. For any open subset $U$ of $X$ containing $Y$, we have $H^{0}(X,\mathcal{L}^{m})\cong H^{0}(U,\mathcal{L}_{U}^{m})\cong H^{0}(X_{/Y},\hat{\mathcal{L}}^{m})$ by the above argument. For any vector bundle $\mathcal{E}$ on $U$, we denote $\mathcal{E}^{\vee}$ as its dual sheaf. Then by coherence of $\mathcal{E}^{\vee}$ and the ampleness of $\mathcal{L}_{U}$, there exist a sufficiently large $m$ that $\mathcal{E}^{\vee}\otimes\mathcal{L}_{U}^{m}$ is generated by global sections, and further sits in the exact sequence
@@
Tensoring $\mathcal{L}^{-m}$ and dualizing, we get exact sequence
@@
Pulling back to $X_{/Y}$ we also have
@@
Therefore we get the commutative diagram with exact rows
@@
The right two vertical maps are isomorphisms thus the left vertical is also an isomorphism. So the condition $\mathrm{Lef}(X,Y)$ holds. ∎
The following is the weakly G3 version of the converse direction of [Har06, Prop. 2.1]. It is much simpler because the weak G3 property is much weaker than the G3 property and as we mentioned in Remark 4.4, the Grothendieck–Lefschetz condition $\mathrm{Lef}(X,Y)$ in the book [Har06] is slightly different from ours.
@@
Let $Y$ be a closed subvariety of an irreducible projective variety $X$ (of dimension $d\geq 2$) that $K(X_{/Y})$ is a field and $\mathcal{L}$ be a line bundle on $X$ that $K(X,\mathcal{L})=K(X)$. If the pair $(X,Y)$ satisfies the condition $\mathrm{Lef}(X,Y)$, then $Y$ is weakly G3 in $X$ with repect to $\mathcal{L}$.
@@
Let $Y$ be a closed subvariety of an irreducible smooth projective variety $X$ of dimension $d\geq 2$, and $\mathcal{L}$ is an ample line bundle on $X$. Then the following are equivalent:
@@
    $Y$ is weakly G3 in $X$ with repect to $\mathcal{L}$ and $Y$ intersects nontrivially with every effective divisor on $X$.

### v2

@@ 1 Introduction
In Kudla’s famous paper [Kud97], he conjectured the modularity of the generating series of special cycles. There are fruitful results in this direction, for example, assuming absolutely convergence of the generating series, Wei Zhang proved the modularity conjecture in the Siegel case in [Zha09] and Yifeng Liu in the unitary case in [Liu12]. To remove the convergence condition, Bruinier and Raum proved the modularity of formal Fourier–Jacobi series in the Siegel case in [BR15] and Jiacheng Xia in some unitary case in [Xia22] by using technical analytic methods. However, recently Bruinier and Raum presented a new approach in this automatic convergence problem in the Siegel case in [BR24], which is closely related to our results on the coherent cohomological dimension of Shimura varieties. Actually, in an ongoing project joint with Wenxuan Qi, Linli Shi, Peihang Wu, Liang Xiao and Yichao Zhang, we have a new approach of the modularity conjecture, where the cohomological dimension appears naturally.
@@
We fix a field $k$ that is algebraically closed and characteristic 0. When we talk about variety, we mean a reduced separated scheme of finite type over $k$, but it does not need to be irreducible. Let $A_{g,\Gamma}$ be the Siegel modular variety of genus $g\geq 2$ and arithmetic level $\Gamma\subset\operatorname{Sp}_{2g}(\mathbb{Q})$ (i.e., $\Gamma$ is commensurable with $\operatorname{Sp}_{2g}(\mathbb{Z})$ and contains a principal congruence subgroup $\Gamma(N)$ for some $N$). Our main result is the following.
@@
As a key application, we deduce that formal Siegel modular forms are automatically convergent, i.e., they are Siegel modular forms. This generalizes the earlier work of Bruinier and Raum [BR15] and [BR24].
@@
To establish Theorem 1.1, we introduce and study the weak G2 and G3 properties with respect to a line bundle, which refine the classical notions of G2 and G3 subvarieties in the sense of Hironaka–Matsumura in [HM68].
@@
Let $A^{\min}_{g,\Gamma}$ denote the minimal compactification, and let $A^{\Sigma}_{g,\Gamma}$ denote the toroidal compactification with respect to the cone decomposition $\Sigma$. Let $\omega$ denote the Hodge line bundle on $A_{g,\Gamma}^{\min}$ or its pullback on $A_{g,\Gamma}^{\Sigma}$. We use $D_{\Gamma}^{\min}$ and $D_{\Gamma}^{\Sigma}$ to denote the corresponding boundary of the compactifications. There is a set-theoretic description of $A_{g,\Gamma}^{\min}$ that
@@
and $D_{\Gamma,l}^{\Sigma}$ as its preimage in $A_{g,\Gamma}^{\Sigma}$.
@@
Acknowledgement. I would like to thank Liang Xiao for asking me to consider this problem. And I want to thank Jan Hendrik Bruinier, Kai-Wen Lan, Wenxuan Qi, Martin Raum, Yichao Tian, Zhiyu Tian, Peihang Wu, Deding Yang, Zhiwei Yun, Wei Zhang, and Yichao Zhang for helpful discussions.
@@ 2 Properties G2 and G3
By [Sta25, Tag 03GE], taking the integral closure commutes with étale base change, so the considered diagram is Cartesian. ∎
@@ 3 Weak G2 and G3 properies with respect to a line bundle
It is obvious that $K(X,\mathcal{L})$ is independent of the choice of the embedding $\mathcal{L}\hookrightarrow K_{X}$.
@@
The part $(1)$ is trivial. For $(2)$, it is trivial that $K(X,\mathcal{L}^{n})\subset K(X,\mathcal{L})$. Conversely, if $s_{0}/s_{1}\in K(\mathcal{L})$ with $s_{0},s_{1}\in H^{0}(X,\mathcal{L}^{m})$, then $s_{0}s_{1}^{n-1},s_{1}^{n}\in H^{0}(X,\mathcal{L}^{nm})$ and therefore $s_{0}/s_{1}=s_{0}s_{1}^{n-1}/s_{1}^{n}$ lies in $K(X,\mathcal{L}^{n})$.
For $(3)$, using $(2)$, we may assume that $\mathcal{L}$ is very ample. This gives a closed immersion $j:X\hookrightarrow\mathbb{P}^{N}$ over $k$ such that $j^{*}\mathcal{O}_{\mathbb{P^{N}}}(1)=\mathcal{L}$, and the coordinates are given by $s_{0},\dots,s_{N}\in H^{0}(X,\mathcal{L})$. Then $K(X)$ can be generated by ratios of $s_{0},\dots,s_{N}$ and therefore is equal to $K(X,\mathcal{L})$. ∎
@@
be the natural quotient morphism. Let $Y$ be a closed subvariety of $X$ such that both $K(X_{/Y})$ and $K(X^{\prime}_{/f^{-1}(Y)})$ are fields. Let $\mathcal{L}$ be a line bundle on $X$ and $\mathcal{L^{\prime}}=f^{*}\mathcal{L}$, such that $K(X,\mathcal{L})=K(X)$ and $K(X^{\prime},\mathcal{L^{\prime}})=K(X^{\prime})$. Then $Y$ is weakly G2 (resp. G3) in $X$ with respect to $\mathcal{L}$ if and only if $f^{-1}(Y)$ is G2 (resp. G3) in $X^{\prime}$ with respect to $\mathcal{L}^{\prime}$.
@@
It is obvious that $G$ is the Galois group of $K(X^{\prime})$ over $K(X)$. The action of $G$ on $X^{\prime}$ induces a natural group action on $H^{0}(X^{\prime}_{/f^{-1}(Y)},\hat{\mathcal{L}^{\prime}})$ and we have
@@
Let $f:X^{\prime}\to X$ be a proper birational morphism of projective varieties, and $Y$ be a closed subvariety of $X$ such that both $K(X_{/Y})$ and $K(X^{\prime}_{/f^{-1}(Y)})$ are fields. Assume further that $X$ is normal. Let $\mathcal{L}$ be a line bundle on $X$ and $\mathcal{L}^{\prime}=f^{*}\mathcal{L}$. Then $Y$ is weakly G2 (resp. G3) in $X$ with respect to $\mathcal{L}$ if and only if $f^{-1}(Y)$ is G2 (resp. G3) in $X^{\prime}$ with respect to $\mathcal{L}^{\prime}$.
@@
It is sufficient to show that $H^{0}(X^{\prime}_{/f^{-1}(Y)},\hat{\mathcal{L}}^{\prime\otimes n})=H^{0}(X_{/Y},\hat{\mathcal{L}}^{\otimes n})$. We show that $f_{*}\mathcal{O}_{X_{/f^{-1}(Y)}^{\prime}}=\mathcal{O}_{X_{/Y}}$ then the above equality holds by the projection formula. By Zariski’s main theorem we have $f_{*}\mathcal{O}_{X^{\prime}}=\mathcal{O}_{X}$, thus we have $f_{*}\mathcal{O}_{X_{/f^{-1}(Y)}^{\prime}}=\mathcal{O}_{X_{/Y}}$ by the flatness of $X_{/Y}\to X$. ∎
@@
Let $g:X_{2}\to X_{1}$ be a proper morphism of normal projective varieties, and let $\mathcal{L}_{1}$ be a line bundle on $X_{1}$ and $\mathcal{L}_{2}=g^{*}\mathcal{L}_{1}$. Let $Y_{1}$ be a closed subvariety in $X_{1}$ and $Y_{2}=f^{-1}(Y_{1})$. Assume that $Y_{1}$ is weakly G2 with respect to $\mathcal{L}_{1}$ and $Y_{2}$ is weakly G2 with respect to $\mathcal{L}_{2}$ in $X_{2}$.
@@
The proof is the same as Proposition 2.12 using the commutative diagram
@@ 4 Cohomological dimension and Grothendieck–Lefschetz condition
In this section, we review the notion of (coherent) cohomological dimension and Grothendieck–Lefschetz condition, and how they are related to the weak G2 and G3 properties.
@@
The following lemma relates $\operatorname{ccd}(X)$ to vanishings of the cohomology groups of twists of negative powers of an ample line bundle.
@@
Let $\mathcal{O}(1)$ be an ample line bundle on $X$. Write $\mathcal{O}(m)$ for its $m$-th power. Then $\operatorname{ccd}(X)\leq r$ if and only if there exist $M\in\mathbb{Z}_{\geq 0}$ and a line bundle $\mathcal{L}$ that $H^{i}(X,\mathcal{L}(-m))=0$ for all $i>r$ and $m\geq M$, where $\mathcal{L}(-m):=\mathcal{L}\otimes\mathcal{O}(-m)$.
@@
It suffices to prove the if part. We prove $\operatorname{ccd}(X)\leq i$ by reverse induction on $i\geq r$. When $i>\dim X$, this is trivial. Suppose that $\operatorname{ccd}(X)\leq i+1$ and $i\geq r$, we prove that $\operatorname{ccd}(X)\leq i$. For any coherent sheaf $\mathcal{F}$ on $X$, we have that $\mathcal{H}(m):=\mathcal{F}\otimes\mathcal{L}^{-1}(m)$ is generated by global sections for sufficiently large $m$ (and $m\geq M$). Then we have an exact sequence
$0\longrightarrow\mathcal{G}\longrightarrow\mathcal{O}_{X}^{\oplus N}\longrightarrow\mathcal{H}(m)\longrightarrow 0$
with a coherent sheaf $\mathcal{G}$, and then an exact sequence
$0\longrightarrow\mathcal{G}\otimes\mathcal{L}(-m)\longrightarrow\mathcal{L}(-m)^{\oplus N}\longrightarrow\mathcal{F}\longrightarrow 0.$
@@
$0=H^{i}(X,\mathcal{L}(-m)^{\oplus N})\longrightarrow H^{i}(X,\mathcal{F})\longrightarrow H^{i+1}(X,\mathcal{G}\otimes\mathcal{L}(-m))=0$
@@
Now, by Serre duality, we have
@@
For local cohomology, we have the exact sequence
@@
and by condition $\operatorname{ccd}(X\backslash Y)\leq d-2$ we have
@@
Combining the duality results above, we have
@@
As $\mathcal{E}_{U}$ is locally free, we have an injective sheaf homomorphism $\mathcal{E}_{U}\hookrightarrow\hat{i}_{*}\hat{i}^{*}\mathcal{E}_{U}$ on $U$, which implies injections
@@
$(3)\Rightarrow(1)$: Since $\mathcal{O}(1)$ is an ample line bundle on $X$, it suffices to prove that $H^{i}(X\backslash Y,\omega_{X}(-m))=0$ when i>d2i>d-2 and $m$ are large enough according to the Lemma 4.2. Again we use the condition and the duality theorems above, we have
$H^{d}_{Y}(X,\omega_{X}(-m))\cong H^{d}(X,\omega_{X}(-m)).$
@@
$H^{d-1}(X,\omega_{X}(-m))\cong H^{1}(X,\mathcal{O}(m))=0$
for $m$ large enough. Now, the expected vanishing follows from the exact sequence
$0\to H^{d-1}(X\backslash Y,\omega_{X}(-m))\to H^{d}_{Y}(X,\omega_{X}(-m))\xrightarrow{\sim}H^{d}(X,\omega_{X}(-m))\to H^{d}(X\backslash Y,\omega_{X}(-m))\to 0.$
The following theorem of Hartshorne and Speiser states a relation between the G3 property and the Grothendieck–Lefschetz condition. See [Bad04, Thm. 10.7] or [Har06, Chap.V, Prop. 2.1].
@@
We need a slightly different version that replaces G3 with weakly G3 with respect to an ample line bundle.
@@
Let $Y$ be a closed subvariety of a projective variety $X$ (of dimension $d\geq 2$) that locally satisfies Serre’s condition $S_{2}$ (e.g. when $X$ is normal). Assume that $Y$ is weakly G3 in $X$ with respect to an ample line bundle $\mathcal{L}$ and $Y$ intersects nontrivially with every effective divisor of $X$. Then $(X,Y)$ satisfies the condition $\mathrm{Lef}(X,Y)$.
@@
The proof is almost the same as the proof of the above theorem. For an integer $m\geq 0$, we fix an embedding $\mathcal{L}^{m}\hookrightarrow K_{X}$, and then induce an embedding $\hat{\mathcal{L}}^{m}:=\hat{i}^{*}\mathcal{L}^{m}\to\mathcal{M}_{X_{/Y}}$ where $\hat{i}:X_{/Y}\to X$ is the natural morphism. We have the commutative diagram
@@
Now we prove that $(X,Y)$ satisfies $\mathrm{Lef}(X,Y)$. For any open subset $U$ of $X$ containing $Y$, we have $H^{0}(X,\mathcal{L}^{m})\cong H^{0}(U,\mathcal{L}_{U}^{m})\cong H^{0}(X_{/Y},\hat{\mathcal{L}}^{m})$ by the above argument. For any vector bundle $\mathcal{E}$ on $U$, we denote $\mathcal{E}^{\vee}$ as its dual sheaf. Then by coherence of $\mathcal{E}^{\vee}$ and ampleness of $\mathcal{L}_{U}$, there exists a sufficiently large $m$ that $\mathcal{E}^{\vee}\otimes\mathcal{L}_{U}^{m}$ is generated by global sections and further sits in the exact sequence
@@
Tensoring $\mathcal{L}^{-m}$ and dualizing, we get an exact sequence
@@
Pulling back to $X_{/Y}$, we also have
@@
Therefore, we get the commutative diagram with exact rows
@@
The two vertical maps on the right are isomorphisms, thus the left vertical is also an isomorphism. So the condition $\mathrm{Lef}(X,Y)$ holds. ∎
The following is the weakly G3 version of the converse direction of [Har06, Prop. 2.1]. It is much simpler because the weak G3 property is much weaker than the G3 property, and as we mentioned in Remark 4.4, the Grothendieck–Lefschetz condition $\mathrm{Lef}(X,Y)$ in the book [Har06] is slightly different from ours.
@@
Let $Y$ be a closed subvariety of an irreducible projective variety $X$ (of dimension $d\geq 2$) that $K(X_{/Y})$ is a field and $\mathcal{L}$ be a line bundle on $X$ that $K(X,\mathcal{L})=K(X)$. If the pair $(X,Y)$ satisfies the condition $\mathrm{Lef}(X,Y)$, then $Y$ is weakly G3 in $X$ with respect to $\mathcal{L}$.
@@
Let $Y$ be a closed subvariety of an irreducible smooth projective variety $X$ of dimension $d\geq 2$, and $\mathcal{L}$ an ample line bundle on $X$. Then the following are equivalent:
@@
    $Y$ is weakly G3 in $X$ with respect to $\mathcal{L}$ and $Y$ intersects nontrivially with every effective divisor on $X$.